# Fair Wasserstein Coresets

### v3

@@ B.1 Downstream Learning using FWC
Overall, the hardness of deriving bounds for the synthetic representatives provided by FWC can be analyzed using the following breakdown, which is adapted from Xu & Strohmer (2023). Consider the given dataset $Z=\{(X_{i},Y_{i},D_{i})\}_{i=1}^{n}$, the synthetic representatives obtained using FWC $\widehat{Z}=\{(\widehat{X}_{j},\widehat{Y}_{j},\widehat{D}_{j})\}_{j=1}^{m}$ and $h\in\mathcal{H}=L^{2}(\mathcal{X}\times\mathcal{Y}\times\mathcal{D})$, the set of measurable square-integrable function in $L^{2}$. If we consider the downstream learning process using FWC samples over the $L^{2}$ space:
@@
The first term corresponds to the loss of information in approximating $Y$ with $\widehat{Y}$ via the FWC approach, and is actually independent of any downstream learning. This condition requires for the first moment (which for the binary $Y$ case is equivalent to the joint distribution) of $Y$ and $\widehat{Y}$ to be as close as possible. Using FWC  this is enforced by minimizing the Wasserstein distance. Indeed, if in definition (1) one restricts to couplings that admit marginal and conditional distributions, then the conditional distributions of $Y|X,D$ and $\widehat{Y}|\widehat{X},\widehat{D}$ are upper bounded in Wasserstein sense by the Wasserstein distance between the joint distribution of $p_{Z}$ and $p_{\widehat{Z}}$ (Kim et al., 2022).
The second terms refers to the training process using FWC samples Firstly, by using the equivalence in Xu & Strohmer (2023), the second term is equivalent to $\inf_{h\in\mathcal{H}}\mathbb{E}_{Y|X,D}\left[\|\widehat{Y}-f(\widehat{X},\widehat{D}\|^{2}_{2}\right]$, which correspond the finding the best $L^{2}$ function to approximate the distribution of $\widehat{Y}$. This fact implies that using FWC samples is indeed mathematically equivalent to the learning task for the original $Y$. However, this second term also highlights the hardness of developing learning bounds, as the FWC synthetic representatives $\widehat{Z}=\{(\widehat{X}_{j},\widehat{Y}_{j},\widehat{D}_{j})\}_{j=1}^{m}$ are not i.i.d., and hence standard bounds are not applicable.
@@ C.1 Runtime Analysis on Synthetic Dataset
This experiment visualizes the speed of our method with respect to different numbers of overall samples $n$, number of samples in the compressed dataset $m$, and the dimensionality of features $p$. We evaluated the performance of the algorithm under the synthetic data with different configurations of $n$, $p$, and $m$. In this experiment, we set compute the fair Wasserstein coreset under the $l_{1}$-norm distance and we use k-means (Lloyd, 1982) to initialize the starting coreset $\widehat{X}^{0}$. We terminate the algorithm when $\widehat{X}^{k}=\widehat{X}^{k-1}$. The time per iteration and total iterations for varying $n$, $p$, and $m$ are shown in the Figures 1 (top left) and 2. We see that increasing the sample size of the original dataset $n$ increases the runtime and number of iterations, while increasing the number of coresets $m$ or dimensionality of the features $p$ reduces the overall numbers of iterations but increases each iteration’s runtime. Additionally, for the setting of Figure 1 (top left), in which we vary the dataset size $n$, Table 2 provides FWC average runtimes from $n=500$ to $n=1,000,000$. We compare FWC runtimes with the runtime at $n=500$ (our lowest dataset size in the experiment) extrapolated (i) linearly, with a factor of 1, (ii) linearly, with a factor of 10 and (iii) quadratically. We can see that the complexity is near linear and less than quadratic with respect to the dataset size $n$, although the rate indeed seem to increase for $n$ at $500,000$ and above, which can be attributed to the increasing number of iterations required to achieve convergence. This is akin to the phenomenon well-known for k-means, for which in larger datasets k-means might take an exponentially large number of iterations to terminate (Vattani, 2009). In practice, a fixed number of overall iterations is set to avoid this case: sklearn sets it to 300300300300333https://scikit-learn.org/stable/modules/generated/sklearn.cluster.KMeans.html, feiss to 25252525444https://faiss.ai/cpp_api/struct/structfaiss_1_1Clustering.html and Matlab to 100100100100555https://www.mathworks.com/help/stats/kmeans.html.
@@ C.2 Real Datasets
We consider the following four real datasets widely used in the fairness literature (Fabris et al., 2022):
  • the Adult dataset (Becker & Kohavi, 1996), which reports demographic data from the 1994 US demographic survey about $\sim 49,000$ individuals. We use all the available features for classification apart from the “fnlwgt” feature, including gender as the protected attribute $D$ and whether the individual salary is more than USD50,0005000050,00050 , 000;
  • the Drug dataset (Fehrman et al., 2017), which contains drug usage history for $1,885$ individuals. Features $X$ include the individual age, country of origin, education and scores on various psychological test. We use the individual gender as the protected variable $D$. The response $Y$ is based on whether the individual has reported to have ever used the drug “cannabis” or not;
  • the Communities and Crime dataset (Redmond, 2009) was put together towards the creation of a software tool for the US police department. The dataset contains socio-economic factors for $\sim 2,000$ communities in the US, along with the proportion of violent crimes in each community. As protected attribute $D$, we include whether the percentage of the black population in the community is above the overall median. For the response $Y$, we use a binary indicator of whether the violent crimes percentage level is above the mean across all communities in the dataset;
  • the German Credit dataset (Hofmann, 1994) reports a set of $1,000$ algorithmic credit decisions from a regional bank in Germany. We use all the available features, including gender as protected attribute $D$ and whether the credit was approved as response $Y$.
@@
  • For Fairlet (Backurs et al., 2019), we use the implementation available at the following GitHub repository: https://github.com/talwagner/fair_clustering/tree/master
  • For IndFair (Chhaya et al., 2022) and K-Median Coresets Bachem et al. (2018), we use the implementation available at the following GitHub repository: https://github.com/jayeshchoudhari/CoresetIndividualFairness/tree/master
  • For k-means (Lloyd, 1982) and k-medoids (Maranzana, 1963; Park & Jun, 2009) we use the implementations available in the Python package Scikit-Learn (Pedregosa et al., 2011)
All computations are run on an Ubuntu machine with 32GB of RAM and 2.50GHz Intel(R) Xeon(R) Platinum 8259CL CPU. For all datasets, we randomly split $75\%$ of the data into training/test set, and change the split during each separate run; the training data are further separated into training and validation with $90/10$ to compute early stopping criteria during training. The downstream classifier used is a one-layer deep multi-layer perceptron (MLP) with $20$ hidden layers, ReLu activation function in the hidden layer and softmax activation function in the final layer. Unless stated otherwise, FWC uses the $L^{1}$ to compute the distance from the original datasets in the optimization problem. For the downstream classifier, we use Adam optimizer (Kingma & Ba, 2014) with a learning rate set to $10^{-3}$, a batch size of $32$, a maximum number of epochs set to $500$ with early stopping evaluated on the separate validation set with a patience of $10$ epochs and both the features $X$ and the protected attribute $D$ are used for training the classifier. Note that due to the size of the Adult dataset, Fairlet coresets Backurs et al. (2019) could not be run due to the RAM memory required exceeding the machine capacity (32GB). Finally, all uncertainties are reported at ±$\pm 1\sigma$ (one standard deviation) in both figures and tables. Uncertainties are computed over a set of 10 runs where the random seed for the algorithm initialization and train/test split was changed, but consistent across all methods (i.e., all methods in the first run were presented the same train/test split across each datasets).
@@
Figure 6 expands the results provided in Figure 1 and shows all the fairness-utility tradeoffs for all methods across the four datasets, both with (right column) and without (left column) using a pre-processing fairness approach (Kamiran & Calders, 2012) (excluding FWC , to which no fairness modification is applied after coresets have been generated). For each method, the coreset size that achieves the best fairness-utility tradeoff is shown (which is not necessarily the coreset with the largest size). FWC  achieves a competitive fairness-utility tradeoff with respect to other competing methods, when using the generated coresets to train a downstream MLP classifier model. FWC consistently reduces disparity in the downstream classification with respect to other approaches, and often maintains the same utility (indicated by the AUC). For completeness, Figure 7 also reports standard deviations for the fairness-utility tradeoff; standard deviations for the Adult and Credit datasets are large due to the MLP classifier becoming trivial (i.e., always returning 00s or $1$s), which yields very low performance but has no demographic disparity (by definition, as all test samples are assigned the same outcome). Tables 5 and 6 report the numbers shown in Figures 6 and 7, including one number to quantify the fairness-utility tradeoff, computed as the Euclidean distance in the Figure from the $(0,1)$ point (which would be a fair classifier with perfect performance). In other words, for the $k$-th method achieving a disparity of $d_{k}$ with uncertainty $\Delta d_{k}$ and performance $a_{k}$ with uncertainty $\Delta a_{k}$, the tradeoff $t_{k}$ and associated uncertainty $\Delta t_{k}$ are quantified as:
@@
Table 6: Demographic disparity (Equation (24)), AUC and fairness-utility tradeoff (Equation (25)) of downstream MLP classifier trained using all fair coresets/clustering methods across the four real datasets. All methods apart from FWC have been corrected for fairness using a preprocessing fairness technique by Kamiran & Calders (2012). The best method across the 3 different coreset sizes is shown, and the best performing method for each metric in each dataset is bolded. Averages and standard deviations taken over 10 runs.
@@
We also evaluate the performance of FWC in reducing the downstream demographic disparity when doing data augmentation, i.e., adding the synthetic representatives to the training data when training a downstream model. We use the data augmentation scheme adopted by Sharma et al. (2020, Section 2.1), which first uses k-means on the original dataset and then sorts the synthetic representatives based on the distance of each synthetic representative to the nearest k-mean centroid with the same combination of protected attribute and outcome $D$ and $Y$. We generate a set of synthetic datasets of size equal to $50\%$ of the original dataset and look at the fairness-utility of a downstream model trained augmented with such synthetic representative in increments of $5\%$ ($20\%$ and $2.5\%$ respectively for the Adult dataset, given the large dataset size). Figures 3 and 4 show the fairness-utility tradeoff of the downstream MLP classifier when doing data augmentation, selecting the best model across various degrees of data augmentation, along with the performance of the baseline MLP classifier with no data augmentation (averages and standard deviations over 10 runs). Table 7 shows the numerical values for the downstream fairness-utility tradeoff, including the tradeoff value computed as in Equation (25). In all datasets the data augmentation with FWC seems to either increase the performance or reduce the demographic disparity, with the only exception being the Drug dataset. Upon further investigation, Figure 5 shows this effect does not appear if the protected attribute $D$ (gender) is not included in the features used to train the downstream MLP classifier. This phenomenon indicates the protected attribute provides a strong predictive effect on the outcome (whether the individual has tried cannabis or not), which might potentially be mediated by unmeasured confounders, i.e., other features regarding the recorded individuals that are not available in the Drug dataset. This would require either in-training or post-processing fairness approaches to be alleviated; see Hort et al. (2023) for a comprehensive review on different potential approaches. Finally, as in Figure 7, the standard deviations for Adult and Credit dataset are large due to the downstream model becoming trivial.
@@
Figure 3: Fairness-utility tradeoff of downstream MLP classifier trained using the original training set augmented with coresets representatives, following the augmentation strategy from Sharma et al. (2020). Each point shows the best model in terms of fairness-utility tradeoff over various degrees of data augmentation, in addition to the baseline model with no augmentation. Means and standard deviations taken over 10 runs.
@@
Figure 4: Fairness-utility tradeoff of downstream MLP classifier trained using the original training set augmented with coresets representatives, following the augmentation strategy from Sharma et al. (2020), including all methods mentioned in Section 7. Each point shows the best model in terms of fairness-utility tradeoff over various degrees of data augmentation, in addition to the baseline model with no augmentation. Averages and standard deviations computed over 10 runs, with the top panel showing just means and the bottom panel combining both means and standard deviations.
@@
To query GPT models, test data with 200 samples is provided with a base parity of 0.5 (similar to Wang et al. 2023) and additionally, examples are provided in the case of the few shot settings. Data is fed in as text. One of the tabular data, for example, is: “A person in 1996 has the following attributes: Age: $21.0$, workclass: Private, education: Some-college, highest education level: $10.0$, marital status: Never-married, occupation: Other-service, relationship: Own-child, race: White, sex: Female, capital gain: $0.0$, capital loss: $0.0$, hours per week: $25.0$, native country: United-States".
@@
In the few shot settings, 16 examples are provided in both cases due to the LLM token limitation; passing fewer examples yields similar results to the ones in Table 1. For FWC, we run a separate coreset generation run (differently from other experiments in Section 7), where we selected $m=16$ and ensured that the positive class ($Y=1$) has an equal number of male and female samples. We also note that while the results from GPT-4 for the zero shot and few shot cases are similar to what was observed by Wang et al. (2023), the accuracies reported for the GPT-3.5 Turbo model appear to be lower in our experiments, which points to a potential difference in the exact backend LLM model used for inference.

### v4

@@ B.1 Downstream Learning using FWC
Overall, the hardness of deriving bounds for the synthetic representatives provided by FWC can be analyzed using the following breakdown, which is adapted from [78]. Consider the given dataset $Z=\{(X_{i},Y_{i},D_{i})\}_{i=1}^{n}$, the synthetic representatives obtained using FWC $\widehat{Z}=\{(\widehat{X}_{j},\widehat{Y}_{j},\widehat{D}_{j})\}_{j=1}^{m}$ and $h\in\mathcal{H}=L^{2}(\mathcal{X}\times\mathcal{Y}\times\mathcal{D})$, the set of measurable square-integrable function in $L^{2}$. If we consider the downstream learning process using FWC samples over the $L^{2}$ space:
@@
The first term corresponds to the loss of information in approximating $Y$ with $\widehat{Y}$ via the FWC approach, and is actually independent of any downstream learning. This condition requires for the first moment (which for the binary $Y$ case is equivalent to the joint distribution) of $Y$ and $\widehat{Y}$ to be as close as possible. Using FWC  this is enforced by minimizing the Wasserstein distance. Indeed, if in definition (1) one restricts to couplings that admit marginal and conditional distributions, then the conditional distributions of $Y|X,D$ and $\widehat{Y}|\widehat{X},\widehat{D}$ are upper bounded in Wasserstein sense by the Wasserstein distance between the joint distribution of $p_{Z}$ and $p_{\widehat{Z}}$ [37].
The second terms refers to the training process using FWC samples Firstly, by using the equivalence in [78], the second term is equivalent to $\inf_{h\in\mathcal{H}}\mathbb{E}_{Y|X,D}\left[\|\widehat{Y}-f(\widehat{X},\widehat{D}\|^{2}_{2}\right]$, which correspond the finding the best $L^{2}$ function to approximate the distribution of $\widehat{Y}$. This fact implies that using FWC samples is indeed mathematically equivalent to the learning task for the original $Y$. However, this second term also highlights the hardness of developing learning bounds, as the FWC synthetic representatives $\widehat{Z}=\{(\widehat{X}_{j},\widehat{Y}_{j},\widehat{D}_{j})\}_{j=1}^{m}$ are not i.i.d., and hence standard bounds are not applicable.
@@ C.1 Runtime Analysis on Synthetic Dataset
This experiment visualizes the speed of our method with respect to different numbers of overall samples $n$, number of samples in the compressed dataset $m$, and the dimensionality of features $p$. We evaluated the performance of the algorithm under the synthetic data with different configurations of $n$, $p$, and $m$. In this experiment, we set compute the fair Wasserstein coreset under the $l_{1}$-norm distance and we use k-means [42] to initialize the starting coreset $\widehat{X}^{0}$. We terminate the algorithm when $\widehat{X}^{k}=\widehat{X}^{k-1}$. The time per iteration and total iterations for varying $n$, $p$, and $m$ are shown in the Figures 1 (top left) and 2. We see that increasing the sample size of the original dataset $n$ increases the runtime and number of iterations, while increasing the number of coresets $m$ or dimensionality of the features $p$ reduces the overall numbers of iterations but increases each iteration’s runtime. Additionally, for the setting of Figure 1 (top left), in which we vary the dataset size $n$, Table 2 provides FWC average runtimes from $n=500$ to $n=1,000,000$. We compare FWC runtimes with the runtime at $n=500$ (our lowest dataset size in the experiment) extrapolated (i) linearly, with a factor of 1, (ii) linearly, with a factor of 10 and (iii) quadratically. We can see that the complexity is near linear and less than quadratic with respect to the dataset size $n$, although the rate indeed seem to increase for $n$ at $500,000$ and above, which can be attributed to the increasing number of iterations required to achieve convergence. This is akin to the phenomenon well-known for k-means, for which in larger datasets k-means might take an exponentially large number of iterations to terminate [73]. In practice, a fixed number of overall iterations is set to avoid this case: sklearn sets it to 300300300300444https://scikit-learn.org/stable/modules/generated/sklearn.cluster.KMeans.html, feiss to 25252525555https://faiss.ai/cpp_api/struct/structfaiss_1_1Clustering.html and Matlab to 100100100100666https://www.mathworks.com/help/stats/kmeans.html.
@@ C.2 Real Datasets
We consider the following four real datasets widely used in the fairness literature [19]:
  • the Adult dataset [7], which reports demographic data from the 1994 US demographic survey about $\sim 49,000$ individuals. We use all the available features for classification apart from the “fnlwgt” feature, including gender as the protected attribute $D$ and whether the individual salary is more than USD50,0005000050,00050 , 000;
  • the Drug dataset [20], which contains drug usage history for $1,885$ individuals. Features $X$ include the individual age, country of origin, education and scores on various psychological test. We use the individual gender as the protected variable $D$. The response $Y$ is based on whether the individual has reported to have ever used the drug “cannabis” or not;
  • the Communities and Crime dataset [64] was put together towards the creation of a software tool for the US police department. The dataset contains socio-economic factors for $\sim 2,000$ communities in the US, along with the proportion of violent crimes in each community. As protected attribute $D$, we include whether the percentage of the black population in the community is above the overall median. For the response $Y$, we use a binary indicator of whether the violent crimes percentage level is above the mean across all communities in the dataset;
  • the German Credit dataset [28] reports a set of $1,000$ algorithmic credit decisions from a regional bank in Germany. We use all the available features, including gender as protected attribute $D$ and whether the credit was approved as response $Y$.
@@
  • For Fairlet [4], we use the implementation available at the following GitHub repository: https://github.com/talwagner/fair_clustering/tree/master
  • For IndFair [12] and K-Median Coresets [3], we use the implementation available at the following GitHub repository: https://github.com/jayeshchoudhari/CoresetIndividualFairness/tree/master
  • For k-means [42] and k-medoids [45, 58] we use the implementations available in the Python package Scikit-Learn [59]
All computations are run on an Ubuntu machine with 32GB of RAM and 2.50GHz Intel(R) Xeon(R) Platinum 8259CL CPU. For all datasets, we randomly split $75\%$ of the data into training/test set, and change the split during each separate run; the training data are further separated into training and validation with $90/10$ to compute early stopping criteria during training. The downstream classifier used is a one-layer deep multi-layer perceptron (MLP) with $20$ hidden layers, ReLu activation function in the hidden layer and softmax activation function in the final layer. Unless stated otherwise, FWC uses the $L^{1}$ to compute the distance from the original datasets in the optimization problem. For the downstream classifier, we use Adam optimizer [38] with a learning rate set to $10^{-3}$, a batch size of $32$, a maximum number of epochs set to $500$ with early stopping evaluated on the separate validation set with a patience of $10$ epochs and both the features $X$ and the protected attribute $D$ are used for training the classifier. Note that due to the size of the Adult dataset, Fairlet coresets [4] could not be run due to the RAM memory required exceeding the machine capacity (32GB). Finally, all uncertainties are reported at ±$\pm 1\sigma$ (one standard deviation) in both figures and tables. Uncertainties are computed over a set of 10 runs where the random seed for the algorithm initialization and train/test split was changed, but consistent across all methods (i.e., all methods in the first run were presented the same train/test split across each datasets).
@@
Figure 6 expands the results provided in Figure 1 and shows all the fairness-utility tradeoffs for all methods across the four datasets, both with (right column) and without (left column) using a pre-processing fairness approach [34] (excluding FWC , to which no fairness modification is applied after coresets have been generated). For each method, the coreset size that achieves the best fairness-utility tradeoff is shown (which is not necessarily the coreset with the largest size). FWC  achieves a competitive fairness-utility tradeoff with respect to other competing methods, when using the generated coresets to train a downstream MLP classifier model. FWC consistently reduces disparity in the downstream classification with respect to other approaches, and often maintains the same utility (indicated by the AUC). For completeness, Figure 7 also reports standard deviations for the fairness-utility tradeoff; standard deviations for the Adult and Credit datasets are large due to the MLP classifier becoming trivial (i.e., always returning 00s or $1$s), which yields very low performance but has no demographic disparity (by definition, as all test samples are assigned the same outcome). Tables 5 and 6 report the numbers shown in Figures 6 and 7, including one number to quantify the fairness-utility tradeoff, computed as the Euclidean distance in the Figure from the $(0,1)$ point (which would be a fair classifier with perfect performance). In other words, for the $k$-th method achieving a disparity of $d_{k}$ with uncertainty $\Delta d_{k}$ and performance $a_{k}$ with uncertainty $\Delta a_{k}$, the tradeoff $t_{k}$ and associated uncertainty $\Delta t_{k}$ are quantified as:
@@
Table 6: Demographic disparity (Equation (25)), AUC and fairness-utility tradeoff (Equation (26)) of downstream MLP classifier trained using all fair coresets/clustering methods across the four real datasets. All methods apart from FWC have been corrected for fairness using a preprocessing fairness technique by [34]. The best method across the 3 different coreset sizes is shown, and the best performing method for each metric in each dataset is bolded. Averages and standard deviations taken over 10 runs.
@@
We also evaluate the performance of FWC in reducing the downstream demographic disparity when doing data augmentation, i.e., adding the synthetic representatives to the training data when training a downstream model. We use the data augmentation scheme adopted by [68, Section 2.1], which first uses k-means on the original dataset and then sorts the synthetic representatives based on the distance of each synthetic representative to the nearest k-mean centroid with the same combination of protected attribute and outcome $D$ and $Y$. We generate a set of synthetic datasets of size equal to $50\%$ of the original dataset and look at the fairness-utility of a downstream model trained augmented with such synthetic representative in increments of $5\%$ ($20\%$ and $2.5\%$ respectively for the Adult dataset, given the large dataset size). Figures 3 and 4 show the fairness-utility tradeoff of the downstream MLP classifier when doing data augmentation, selecting the best model across various degrees of data augmentation, along with the performance of the baseline MLP classifier with no data augmentation (averages and standard deviations over 10 runs). Table 7 shows the numerical values for the downstream fairness-utility tradeoff, including the tradeoff value computed as in Equation (26). In all datasets the data augmentation with FWC seems to either increase the performance or reduce the demographic disparity, with the only exception being the Drug dataset. Upon further investigation, Figure 5 shows this effect does not appear if the protected attribute $D$ (gender) is not included in the features used to train the downstream MLP classifier. This phenomenon indicates the protected attribute provides a strong predictive effect on the outcome (whether the individual has tried cannabis or not), which might potentially be mediated by unmeasured confounders, i.e., other features regarding the recorded individuals that are not available in the Drug dataset. This would require either in-training or post-processing fairness approaches to be alleviated; see [29] for a comprehensive review on different potential approaches. Finally, as in Figure 7, the standard deviations for Adult and Credit dataset are large due to the downstream model becoming trivial.
@@
Figure 3: Fairness-utility tradeoff of downstream MLP classifier trained using the original training set augmented with coresets representatives, following the augmentation strategy from [68]. Each point shows the best model in terms of fairness-utility tradeoff over various degrees of data augmentation, in addition to the baseline model with no augmentation. Means and standard deviations taken over 10 runs, with the computed Pareto frontier indicated by the dashed red line.
@@
Figure 4: Fairness-utility tradeoff of downstream MLP classifier trained using the original training set augmented with coresets representatives, following the augmentation strategy from [68], including all methods mentioned in Section 7. Each point shows the best model in terms of fairness-utility tradeoff over various degrees of data augmentation, in addition to the baseline model with no augmentation. Averages and standard deviations computed over 10 runs, with the top panel showing just means and the bottom panel combining both means and standard deviations, with the computed Pareto frontier indicated by the dashed red line.
@@
To query GPT models, test data with 200 samples is provided with a base parity of 0.5 (similar to [76]) and additionally, examples are provided in the case of the few shot settings. Data is fed in as text. One of the tabular data, for example, is: “A person in 1996 has the following attributes: Age: $21.0$, workclass: Private, education: Some-college, highest education level: $10.0$, marital status: Never-married, occupation: Other-service, relationship: Own-child, race: White, sex: Female, capital gain: $0.0$, capital loss: $0.0$, hours per week: $25.0$, native country: United-States".
@@
In the few shot settings, 16 examples are provided in both cases due to the LLM token limitation; passing fewer examples yields similar results to the ones in Table 1. For FWC, we run a separate coreset generation run (differently from other experiments in Section 7), where we selected $m=16$ and ensured that the positive class ($Y=1$) has an equal number of male and female samples. We also note that while the results from GPT-4 for the zero shot and few shot cases are similar to what was observed by [76], the accuracies reported for the GPT-3.5 Turbo model appear to be lower in our experiments, which points to a potential difference in the exact backend LLM model used for inference.
Appendix D Limitations of FWC
Coreset support and non-convex feature spaces FWC representative do not need to be within the original $n$ samples, but they do need to share the same support. As shown in Section 4.2, solving line 5 in Algorithm 1 in cases 1. and 2. means the synthetic representatives could fall outside the original dataset. In case 3., the solution has to be selected from the data points already existing in the original dataset (akin to k-medoids). In general, non-convex feature spaces $\mathcal{X}$ might represent a challenge, as representatives might be generated in zero-density regions (a simple example could be a dataset distributed as a hollow circle or two moons). However, this criticism is also more generally applicable to the existing fair coresets/clustering literature, as well as the k-means algorithms, for which specific adjustments have been developed [63] and could be indeed extended to FWC .
Limiting the total number of iterations FWC is not of polynomial time complexity and the total number of iterations might grow faster than linear time when the dataset size is very large, as shown in the synthetic data experiment in Section C.1 and Table 2. This phenomenon is shared with k-means, which is also not of polynomial time complexity and is known to potentially take an exponentially large number of iterations to terminate [73]. As mentioned in Section C.1, a common practice for clustering is set a fixed number of maximum iterations, after which the algorithm is stopped.
Computational bottlenecks The main complexity term for FWC is $\mathcal{O}(mn)$, which comes from establishing the cost matrix in the beginning of the solution of problem (8). This complexity is comparable with what occurs in Lloyd’s algorithm for k-means and k-medians. This might be problematic if the cost matrix is too large to be stored directly in memory. In practice, we do not actually need to store the entire matrix, as we only need to compute the largest component for each row of $C$ for solving problem (8) (see Lemma A.1), so one could further improve the cost of storing the cost matrix. Another option would also be leverage the same approaches used for k-means such as, e.g., cost matrix sketching [79]. Finally, FWC would also benefit from GPU implementations akin to k-means and k-medians, which would substantially accelerate the runtime speed of FWC .
Connection between $\epsilon$ and downstream learning In our definition of demographic parity in Equation  3, the hyper-parameter $\epsilon$ effectively controls how different the outcome rates across sensitive feature groups $D$ of the weighted coreset distribution $p_{\widehat{Z},\theta}$ can be from the overall outcome rates in the original dataset $p_{Y_{T}}$. In our experiments (Section 7), we empirically show that limiting the fairness violation in the coresets results in a fairer downstream model. However, when training a downstream model using FWC we induce a distribution shift between the train set and the test set, as the coresets distribution is never identical to the original dataset distribution. Although we have provided some results on the generalization properties of FWC (Proposition 5.5) as well as some intuition about developing downstream learning bounds for FWC in non-i.i.d. settings (Section B.1), theoretically characterize the connection between the fairness violation parameter $\epsilon$ remains an open question. In essence, the analysis is challenging as the induced distribution shift is dependent on the biases in the original dataset distribution, the coreset size $m$, the metric chosen for the cost matrix and, ultimately, the fairness violation parameter $\epsilon$. For this reason, although we have shown that restricting the fairness violation improves downstream models fairness, an explicit characterization of the downstream effects of $\epsilon$, as well as other hyper-parameters, would require significant further work, beyond the scope of this paper.
Implications for other fairness measures As highlighted by [5, Chapter 3], fairness notions in classification settings can be categorized into notions of independence, separation, and sufficiency. Demographic parity falls in the class of the independence notion, and hence, other measures of fairness that are closely related, e.g., disparate impact, would also improve when optimizing for demographic parity. However, other notions of fairness such as separation or sufficiency may not simultaneously be satisfied [5]. As FWC targets demographic parity, it cannot guarantee an improvement in these other measures. To test this, we compute the equalized odds, which falls under the notion of separation, for the downstream classifier in Section 7 and check the performance of FWC compared to the other approaches. Table 8 indicates the datasets in which FWC is part of the Pareto frontier for both demographic parity and equalized odds. When considering equalized odds, FWC is not a part of the Pareto frontier for the Drug dataset, and more generally, FWC performance is not as competitive. This is in contrast to demographic parity: in Figure 1, FWC sits on the Pareto frontier across all datasets for fairness-performance trade-off in downstream classification.
Table 8: Presence on the Pareto frontier for FWC across different fairness violation hyper-parameter values ($\epsilon=\{0.01,0.05,0.1\}$), for both demographic parity (left) and equalized odds (right) in the downstream learning settings of Section 7. As equalized odds is not an independence notion of fairness as demographic parity, constraints on demographic parity do not guarantee an improvement in equalized odds, resulting in FWC not performing as well for downstream performance-fairness tradeoff when using equalized odds.
Pareto Frontier, Demographic Parity Pareto Frontier, Equalized Odds
Dataset FWC ($\epsilon=0.01$) FWC ($\epsilon=0.05$) FWC ($\epsilon=0.1$) FWC ($\epsilon=0.01$) FWC ($\epsilon=0.05$) FWC ($\epsilon=0.1$)
Adult $\checkmark$ $\checkmark$ $\checkmark$
Drug $\checkmark$ $\checkmark$ $\checkmark$
Crime $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Credit $\checkmark$ $\checkmark$ $\checkmark$
Appendix E Broader Impact
Our work presents a novel approach to obtain coresets (synthetic representative samples) of a given dataset while reducing biases and disparities in subgroups of the given dataset. As other approaches in the field of algorithmic fairness, our efforts may help populations that would otherwise face disadvantages from a model or decision process. Importantly, our approach refrains from exploiting biases inherent in the data itself; rather, it seeks to mitigate biases in data-driven decision systems. It is crucial to note that our method does not claim to address all sources or types of bias. In addition, while our tools enable a malicious modeler to manipulate algorithmic fairness methods to amplify disparities instead of reducing them, for instance, by reversing the fairness constraint (replacing $\leq$ with $\geq$), the unfairness of a trained model can be detected by assessing it over a separate test set from the original dataset.